OpenStudy (rockinhood):

I will Fan you, Give you a medal and write a testimony if you help me answer this question.

OpenStudy (rockinhood):

Ralph has been awarded some money in a settlement. He has the option to take a lump sum payment of $425,000 or get paid an annuity of $2,000 per month for the next 25 years. Which is the better deal for Ralph, and by how much, assuming the growth rate of the economy is 5.15% per year? Lump Sum: by $91,772.64 Lump Sum: by $34,251.30 Annuity: by $91,772.64 Annuity: by $34,251.30

OpenStudy (rockinhood):

For Lump sum it's FV PV = ________ (1 + i)^nt

OpenStudy (rockinhood):

Annuity is http://prntscr.com/apd6lc

OpenStudy (rockinhood):

@FaiqRaees @welshfella @sweetburger @imqwerty

OpenStudy (anonymous):

present value of the annuity PV = PMT(1- (1+r)^-n)/r with r & n adjusted for periodicity = 2000(1- (1+5.15/1200)^-300)/(5.15/1200) = $337,062 take the lump sum its PV is better by (425000 - 337062) = $87,938 <------ note: ------- it has been assumed that payments are at the end of each month, ie start 1 month after winning the prize. if they start immediately, the PV of the annuity will be $338,509 and the benefit of lump sum will be $86,491

OpenStudy (rockinhood):

I don't know if that actually helps because I already saw this exact thing word for word on Yahoo Answers....

OpenStudy (rockinhood):

@sweetburger ??

OpenStudy (rockinhood):

These are the answers so your logic doesn't match up Lump Sum: by $91,772.64 Lump Sum: by $34,251.30 Annuity: by $91,772.64 Annuity: by $34,251.30